OpenStudy (anonymous):

Solve for x: 2 over 5 plus 3 over 5 x equals quantity x plus 5 over 10 x = –3 x = 2 x = –3 and x = 2 No Solution

OpenStudy (anonymous):

im thinking the answer is either -3 or no solution but im not sure

OpenStudy (anonymous):

Does it look like this @marinesrock? \[Solve \frac{ 2 }{ 5 } + \frac{ 3 }{ 5x } = x + \frac{ 5 }{ 10 } for x\] ???

OpenStudy (anonymous):

yea that is it.i need to solve for x

OpenStudy (anonymous):

Well the solution I got was \[x =\frac{ -1 }{ 4 } \approx -0.25\]

OpenStudy (anonymous):

ok so it is no solution.thank you

OpenStudy (anonymous):

You're welcome!